Help
CH 15 HW Problem 15.67 Problem 15.67 The 12 g head of a bobble-head doll oscillates in SHM at a frequency of 5.9 Hz Resources previous l 9 of 13 l nex Correct Significant Figures Feedback: Your ans 16.49092 N was either rounded differently or used a different number of significant figures than required for this pa Part B Suppose the head is pushed 2.0 cm against the spring, then released. The amplitude of the heads oscillations decreases to 0.5 cm in 4.0 s. What is the heads damping constant? Express your answer using two significant figures. kg/s Submit My Answers Give Up Provide Feedback Continue
Answer #1

w = 2 pi f = sqrt [k / m]

k = m * (2 pi f)2 = 12 * 10-3 * (2 pi * 5.9)2= 16.49 N/m

v = w A = xm * sqrt [k / m] = 0.02 * sqrt [16.49 / 12 * 10-3] = 0.741 m/s

x(t) = xm * e-bt/2m

0.005 = 0.02 * e-bt/2m

e-bt/2m = 0.25

-bt/2m = ln (0.25) = -1.386

-b(4.0 s) / (2 * 0.012) =  -1.386

b = 8.316 * 10-3 kg/s